OpenStudy (anonymous):

Is My Answer Correct? The height of a trapezoid is 13 inches, one of the bases measures 22 inches. If the area of the trapezoid is 318.5 square inches, find the length of the other base. I think it's 27, but i'm not sure.

OpenStudy (anonymous):

1/2(22+x)*13=318.5 (11+1/2x)*13=318.5 143+7.5x=318.5 7.5x=175.5 x=175.5/7.5 23.4

OpenStudy (anonymous):

You missed from the 3rd line, 6.5 instead of 7.5 143 + 6.5x = 318.5 6.5x = 175.5 x = 27 It is much easier if you multiply both sides by 2 from beginning. (22 + x) 13 = 637 You then divide both side by 13 to obtain 22 + x = 49 Fianlly subtract both sides by 22 x = 27